# G2D2: Gradient-guided Discrete Diffusion for image inverse problem solving

### v1

@@ C.6 Impact of the Forget Coefficient
We present additional qualitative results of G2D2 and comparison methods. Figures 8 through 11 showcase the results for super-resolution (SR) and Gaussian blur (GB) tasks on ImageNet and FFHQ datasets.
@@
C.8 Additional qualitative results of G2D2 with Markov noise process
To compare G2D2 and G2D2 with Markov noise process, we present their respective qualitative results in Figures 12 and 13. The latter approach does not include re-masking operations in its sampling process, which means that once a token becomes unmasked, it cannot be modified in subsequent iterations. The unnatural artifacts observed in the resulting images are likely attributable to this limitation. This observation underscores the validity of adopting the star-shaped noise process in our proposed method.
@@ C.8 Additional qualitative results of G2D2 with Markov noise process
C.9 Inverse problems on motion data
We develop G2D2 based on the official implementation of MMM (Pinyoanuntapong et al., 2024)  [^12] https://github.com/exitudio/MMM. This method learns a masked generative model on the discrete latent space obtained by a motion tokenizer trained on the VQVAE framework (Van Den Oord et al., 2017). G2D2 uses the provided pre-trained model as a prior distribution.
@@ C.9 Inverse problems on motion data
In our experiments, we use two types of paths (forward: a path moving forward at a constant speed, and zigzag: a path moving forward while zigzagging) and two prompts: ”A person walks with his hands up” and ”A person does a cart wheel.” We also perform unconditional generation. The qualitative results are shown in Figure 14.
We conduct experiments with a total of $T=25$ time steps. For hyperparameters, we set the number of iterations for optimization to 20 and the Gumbel-Softmax temperature to 1.0. The forget coefficient is set to $0.7$. We adopt the dynamic learning rate scheduling described in the Appendix C.2. The base Adam learning rate $l_{\text{Adam, base}}$ is set to 0.3, and the KL divergence weight $\eta_{\text{KL}}$ is set to 0.05. Additionally, we set $\lambda_{\text{KL, schedule}}$ and $\lambda_{\text{lr, schedule}}$ to 0.0 and 1.0, respectively.
Figure 14: Qualitative results of motion inverse problem solving.

### v2

@@ E.9 Failure modes of G2D2
We present additional qualitative results of G2D2 and comparison methods. Figures 10 through 13 showcase the results for super-resolution and Gaussian deblurring tasks on ImageNet and FFHQ datasets.
@@
E.12 Additional qualitative results of G2D2 with Markov noise process
To compare G2D2 and G2D2 with Markov noise process, we present their respective qualitative results in Figures 14 and 15. The latter approach does not include re-masking operations in its sampling process, which means that once a token becomes unmasked, it cannot be modified in subsequent iterations. The unnatural artifacts observed in the resulting images are likely attributable to this limitation. This observation underscores the validity of adopting the star-shaped noise process in our proposed method.
@@ E.12 Additional qualitative results of G2D2 with Markov noise process
E.13 Inverse problems on motion data
We develop G2D2 based on the official implementation of MMM (Pinyoanuntapong et al., 2024)  [^14] https://github.com/exitudio/MMM. This method learns a masked generative model on the discrete latent space obtained by a motion tokenizer trained on the VQ-VAE framework (van den Oord et al., 2017). G2D2 uses the provided pre-trained model as a prior distribution.
@@ E.13 Inverse problems on motion data
We conduct experiments with a total of $T=25$ time steps. For hyperparameters, we set the number of iterations for optimization to 20 and the Gumbel-Softmax temperature to 1.0. The forget coefficient is set to $0.7$. We adopt the dynamic learning rate scheduling described in the Appendix E.2. The base Adam learning rate $l_{\text{Adam, base}}$ is set to 0.3, and the KL divergence weight $\eta_{\text{KL}}$ is set to 0.05. Additionally, we set $\lambda_{\text{KL, schedule}}$ and $\lambda_{\text{lr, schedule}}$ to 0.0 and 1.0, respectively.